OpenStudy (anonymous):

There were 170 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $490. How many Juniors bought tickets? 70 80 90 100

OpenStudy (anonymous):

i think it was like 80

OpenStudy (anonymous):

two ways to solve you can write \[x+y=170\] that is, the total of the two types of tickets sold is 170 and \[2x+3.5y=490\] i.e. the total amount of money was 490 then solve this

OpenStudy (anonymous):

i got the answer to be 80

OpenStudy (anonymous):

or you can put \(x\) equal the number of $2 tickets so \(170-x\) is the number of $3.50 tickets and solve \[2x+3.5(170-x)=490\] fpr \(x\)

OpenStudy (anonymous):

70

OpenStudy (anonymous):

no i do not think it is 80 yes, 70

OpenStudy (anonymous):

i got 70

OpenStudy (anonymous):

good

OpenStudy (anonymous):

Thanks!!